OpenStudy (anonymous):

Alex designed the model of a cylindrical water tank to be installed in a park. The model has a radius of 6 cm and a height of 13 cm. The dimensions of the actual water tanks that will be installed in the park are 6 times larger than the model. The total surface area of the actual water tanks, rounded to the nearest whole number, is _________ cm². Use twentytwosevenths_large for pi.

OpenStudy (anonymous):

by six times larger i assume that it means the radius will be 36 cm and the height will be 78 cm, right?

OpenStudy (anonymous):

area of bottom is \(\pi r^2=\pi\times 36^2=1296\pi\) and the area of the top is the same, so for the top and bottom together the total area will be \[2\times 1296\pi=2592\pi\] now we need the area of the cylinder

OpenStudy (anonymous):

height is 78, base is \(2\pi r =2\pi \times 36=72\pi\) so area of cylinder part is \[78\times 72\pi=5616\pi\] and so total surface area is \[2596\pi+5616\pi=8212\pi\] check my arithmetic

OpenStudy (anonymous):

Thank you very much for your help, this is much clear to me now :)